# Estimating the link budget of satellite-based Quantum Key Distribution (QKD) for uplink transmission through the atmosphere

### v1

@@ 1 Introduction
On the contrary, the uplink configuration lacks constraints regarding photon source quantity or weight. It adds more flexibility to any change of the photon source required in the future, which is possible with the source on the ground. Also, it will allow for the inclusion of supplementary optical components, such as those required for error correction and characterization. However, in the case of uplink, the beam encounters the atmosphere in the initial stage of its propagation. The effect of the atmosphere on attenuating the link and reducing the key rate is unavoidable. Hence, atmospheric simulation plays an important role in modeling uplink compared to downlink-based QC.
@@
The Canary Island experiments [12, 13, 14] present experimental results, indicating complete losses where turbulence accounts for 18 dB, atmospheric transmittance contributes 15 dB, and the remaining losses are attributed to optical factors, resulting in a total loss of 40 dB. The initial step involves simulating these outcomes with all the parameters to validate the loss values. The procedure’s specifics are outlined in sections A and A.1 of the appendix. This horizontal free-space communication spanning 144 kilometers exhibits losses that are equivalent to those experienced in the LEO. In fact, simulations indicate that the anticipated link transmittance from an LEO satellite will be quite similar [21]. In contrast, the Canadian study  [20], contains simulation findings of satellite loss in LEO orbit at various locations across Canada. To demonstrate that our simulation is accurate, we extended it to the LEO orbit and compared the findings with those from Canada. The detailed verification is given in A.2 of the Appendix. After comparing our simulation strategy with experimental results [12, 13, 14] and simulations conducted in other geographical regions  [20], we proceed to apply it in detail to diverse scenarios within India. Specifically, our focus is on three distinct regions. for three well-known observatories in the near-IR and optical bands. One of the observatories is the Indian Astronomical Observatory (IAO) in Hanle which is the highest astronomical observatory in India at 4500 meters altitude in the Himalayan region  [22]. Situated in the Shivalik mountain range, the Aryabhatta Research Institute of Observational Sciences (ARIES) in Nainital serves as the second observatory  [23]. ARIES is 1951 meters above sea level, and during the observation, the monsoon plays a crucial role. The third observatory is near Mount- Abu, which is located close to the Thar Desert at 1680 meters above sea level, where there is almost no rain for most of the year. The varied atmospheric conditions make India a fascinating study for varied geographical regions and their applicability to space-based QKD.
@@ 2 Link Budget analysis
where $G_{t}$, which is the gain of the transmitter, $G_{r}$ is the receiver antenna gain and $L_{r}$ is the Free space propagation loss given by,
@@
1 Tx gain ($G_{t}$) dB 103.5
2 Tx Beam divergence ($2\Theta_{B}$) $\mu rad$ 19.85
@@
4 Path loss ($L_{r}$) dB -258
5 Atmospheric attenuation ($\eta_{atm}$)=0.325 dB -4.87
7 RX gain ($G_{r}$) dB 121.35
@@ 4 Path loss ($L_{r}$) dB -258
9 RX pointing loss $\mu rad$ 2
10 Total loss dB 44.26
Table 1: Link budget calculation
@@ 2.1 Transmitter:
A telescope converts the source’s quantum beam into a signal that is sent to the receiving satellite. The beam mainly depends upon the aperture size of the telescope and divergence. A smaller beam size leads to greater losses at the receiver due to its higher divergence. On the other hand, while a larger beam size exhibits less divergence, it becomes more susceptible to turbulence. Therefore, achieving an optimal beam size is crucial for improved performance. The gain of the transmitter beam [26], $G_{t}$ is given in (3). ${\Theta_{B}}$ is related to wavelength $\lambda$ and beam size $W_{0}$ as $\Theta_{B}=\frac{\lambda}{\pi W_{0}}$.
@@ 2.2 Path loss:
Given its proximity to Earth, LEO experiences minimal diffraction loss, making it particularly advantageous for quantum key distribution (QKD) experiments. However, orbits with slower speeds, such as the Middle Earth Orbit (MEO) or GEO, offer the benefit of maintaining a continuous link for extended periods, allowing for prolonged QKD operations. Nonetheless, it’s worth noting that MEO and GEO orbits come with their own challenges, including higher radiation levels and increased propagation loss. Free space propagation loss [30] in dB is $\log_{10}L_{r}$ where $L_{r}$ is given in (3).
@@ 2.2.1 Atmospheric loss:
$\eta_{atte}$ is due to the static components of the atmosphere and $\eta_{tur}$ is due to the turbulence i.e dynamic components.
The Static components constitute various types of gases, water vapors, and dust particles that constitute the static atmosphere. These atmospheric elements absorb energy from the photons and this phenomenon is known as absorption. It depends on the wavelength. Photon collisions with air particles cause scattering  [31]. The dynamic components are discussed in the next 2.2.2 subsection.
@@
Table 2: Atmospheric parameter consideration for MODTRAN plot for IAO Hanle
Figure 1: Atmospheric transmittance of IAO Hanle
@@ 2.2.2 Turbulence
$C^{2}_{n}(h)=0.00594\bigg{(}\frac{v}{27}\bigg{)}^{2}(10^{-5}h)e^{-10^{-3}h}+2.7\times 10^{-16}e^{-\frac{2h}{3\times 10^{3}}}+A_{0}\ e^{-10^{-2}h}\,.$ (7)
@@
When there is no turbulence present, the parameter $r_{0}$ represents the diameter of an equivalent aperture where the resolution of a telescope is approximately diffraction-limited. A larger value of $r_{0}$ implies that turbulence has a lesser impact on the propagation of the beam and vice versa. Since $r_{0}$ varies with the wavelength according to $\lambda^{6/5}$, longer operating wavelengths result in larger values of the $r_{0}$. This suggests that at longer wavelengths, the effect of turbulence on the wavefront is less severe.
One more way to express the strength of turbulence is $\frac{D}{r_{0}}$. This is a ratio of the diameter of the beam from the telescope aperture to the atmosphere’s coherence length. The higher the value, the more turbulated the beam. As the value for $\frac{D}{r_{0}}$ increases higher order wavefront distortion will accumulate  [37]. Hence objective is to have minimum $\frac{D}{r_{0}}$. But its value shouldn’t be $<1$. Though no wavefront distortion will be there but optical beam will go through a beam wander phenomenon which makes it difficult to detect the beam at the receiver.
@@ 3 Link budget calculation for beacon beams
To establish and sustain a free-space optical link, it is essential for the spacecraft’s altitude determination and control system to ensure precise pointing accuracy toward the optical ground station. Conversely, a laser beacon tracking system has the capability to furnish accurate ground-based altitude information, thus facilitating communication from the satellite. A laser beacon is transmitted from the optical ground station with a different wavelength from the quantum signal beam. The satellite will also transmit a beacon beam to get the position of the ground station. For the uplink, we use 532 nm, and for the downlink beacon 1550 nm wavelength. The link budget calculation for the beacon beam is also given in Table 3 for a ground telescope of 15cm aperture and a satellite telescope of 30cm aperture  [43]. The calculation for the optimum size of 15cm is given in section A.4 of the Appendix.
Parameter unit Uplink Downlink
Tx power ($P_{t}$) W 1.0 1.0
Tx Beam waist($1/e^{2}$) mm 150.0 300.0
Wavelength ($\lambda$) nm 532 1550
Tx Beam divergence ($2\Theta_{B}$) $\mu rad$ 500 500
Tx optics loss ($\eta_{t}$) dB -2.20 -2.20
Tx gain ($G_{t}$) dB 71.23 69.06
Propagation distance ($L$) km 500 500
Path loss ($L_{r}$) dB -258 -258
Atmospheric absorption ($\tau_{atm}$)=0.7/0.8 dB -1.55 -.97
Turbulence ($\tau_{tur}$)=0.5 dB -3 -3
RX antenna diameter cm 30.0 15.0
RX gain ($G_{r}$) dB 125.0 118.65
RX optics loss ($\eta_{r}$) dB -2.2 -2.2
total loss dB 72.36 74.56
Table 3: Link budget for uplink 532nm and downlink 1550 nm beam.
Table 1 gives the link budget for the signal beam, Table 3 gives the link budget for the uplink and downlink beacon beams. The atmospheric transmittance for beacon beams is taken from Table 2. The link budget analysis for IAO Hanle with Aries Nanital and Mount-Abu is given and compared in Table 4. Data regarding the atmospheric conditions at all of these locations has been acquired from the GIOVANNI database for the year 2015 Jan to Dec. For a comprehensive understanding of the parameters employed to compute the link budget for these two observatories, please refer to Table A4 within the Appendix section. The 2nd and 3rd column gives the loss for the beacon beam at 532 nm (uplink) and 1550 nm (downlink), whereas the last column for the signal beam is at 810 nm (uplink). For the downlink beacon due to the smaller size of the ground telescope, the loss is approximately 2 dB more compared to the uplink.
Location uplink loss(dB) downlink loss(dB) signal loss(dB)
IAO Hanle 72.36 74.56 44.28
ARIES Nainital 74.17 76.48 48.8
Mount-Abu 75.85 78.16 47.50
Table 4: Link budget analysis for 3 different locations in India
@@
When the zenith angle is zero degrees, the loss is indicated in Table 4 for the three separate places in India. The air mass will increase when the satellite moves at a greater zenith angle, which is the cause of an increase in loss in dB, as depicted in Fig. 2. The plot is given for one of the locations IAO Hanle.
Figure 2: Atmospheric transmission at different angles
@@
4 Quantum signal and final key rate
The primary objective is to facilitate secure key exchange while minimizing the potential leakage of information to the third party (eavesdropper). Additionally, examining the Quantum Bit Error Rate (QBER) enables users to identify any potential eavesdropping on the quantum channel. The QBER in a QKD system is essentially the ratio of incorrect counts to the overall number of received counts. Errors in a free-space Quantum Key Distribution (QKD) system generally originate from two distinct sources. The first source involves a base error rate that arises due to imperfections in the optical components of the system. The second source is detector noise, which comprises two components: detector dark count and background radiation. Therefore, the error rate $E$ in a free-space QKD system can be defined as follows: [44]
$E=k+\frac{2Bt}{\mu T\eta}\,,$ (20)
  • $k$ is the error rate contribution from the optical system.
  • $B$ is the background count rate per second.
  • $t$ is the width of the detector timing gate.
  • $\mu$ is the average photon number.
  • T is the transmission of the channel.
  • $\eta$ is the detector efficiency.
On the receiving end of the quantum channel, Bob receives a portion of Alice’s transmission. Some photons are absent due to losses occurring within the channel. When Alice and Bob possess nearly identical lists, errors will arise inevitably due to sources of noise such as the transmission medium, imperfections in the physical apparatus etc. The error correction scheme performs optimally when dealing with error rates $E<0.07$(assuming a 0.02 base error rate due to optical element imperfections). As a result, we can observe that the lowest atmospheric transmission corresponds directly to the dark count rate, assuming a timing gate of t = 1 ns.
$T>\frac{4\times 10^{-8}B}{\mu\eta}\,,$ (21)
The error rate remains unaffected by the pulse repetition rate. However, as the level of loss increases, the key rate $K$ will decrease. Since $K=R\mu T\eta$ where R is the repetition rate of the source.

### v2

@@ 1 Introduction
On the contrary, the uplink configuration lacks constraints regarding photon source quantity or weight. It adds more flexibility to any change of the photon source required in the future, which is possible with the source on the ground. Also, it will allow for the inclusion of supplementary optical components, such as those required for error correction and characterization. However, in the case of uplink, the beam encounters the atmosphere in the initial stage of its propagation. The effect of the atmosphere on attenuating the link and reducing the key rate is unavoidable. Hence, atmospheric simulation plays an important role in modelling uplink compared to downlink-based QC.
@@
The Canary Island experiments [12, 13, 14] present experimental results, indicating complete losses where turbulence accounts for 18 dB, atmospheric transmittance contributes 15 dB, and the remaining losses are attributed to optical factors, resulting in a total loss of 40 dB. The initial step involves simulating these outcomes with all the parameters to validate the loss values. The procedure’s specifics are outlined in sections A and A.1 of the appendix. This horizontal free-space communication spanning 144 kilometers exhibits losses that are equivalent to those experienced in the LEO. In fact, simulations indicate that the anticipated link transmittance from an LEO satellite will be quite similar [21]. In contrast, the Canadian study  [20], contains simulation findings of satellite loss in LEO orbit at various locations across Canada. To demonstrate that our simulation is accurate, we extended it to the LEO orbit and compared the findings with those from Canada. The detailed verification is given in A.2 of the Appendix. After comparing our simulation strategy with experimental results [12, 13, 14] and simulations conducted in other geographical regions  [20], we proceed to apply it in detail to diverse scenarios within India. Specifically, our focus is on three distinct regions for three well-known observatories in the near-IR and optical bands. One of the observatories is the Indian Astronomical Observatory (IAO) in Hanle which is the highest astronomical observatory in India at 4500 meters altitude in the Himalayan region  [22]. Situated in the Shivalik mountain range, the Aryabhatta Research Institute of Observational Sciences (ARIES) in Nainital serves as the second observatory  [23]. ARIES is 1951 meters above sea level, and during the observation, the monsoon plays a crucial role. The third observatory is near Mount- Abu, which is located close to the Thar Desert at 1680 meters above sea level, where there is almost no rain for most of the year. The varied atmospheric conditions make India a fascinating study for varied geographical regions and their applicability to space-based QKD.
@@ 2 Link Budget analysis
where $G_{t}$, which is the gain of the transmitter, $G_{r}$ is the receiver telescope gain and $L_{r}$ is the Free space propagation loss given by,
@@
1 Tx gain ($G_{t}$) dB 109.03
2 Tx Beam divergence ($2\Theta_{B}$) $\mu rad$ 20
@@
4 Path loss ($L_{r}$) dB -257.79
5 Atmospheric attenuation ($\eta_{atm}$)=0.651 dB -1.84
6 Beam Wander loss $(L_{BW})$ dB -0.40
7 RX gain ($G_{r}$) dB 121.32
@@
9 RX pointing loss dB -1.83
10 Total loss dB 35.91
Table 1: Link budget calculation for the IAO Hanle observatory. Row 1, 2, and 3 represent parameters for the transmitter telescope. Row 4, 5 and 6 represent propagation losses. Row 7, 8, and 9 represent parameters for the receiver telescope.
@@ 2.1 Transmitter:
A telescope converts the source’s quantum beam into a signal that is sent to the receiving satellite. The beam mainly depends upon the aperture size of the telescope and divergence. A smaller beam size leads to greater losses at the receiver due to its higher divergence. On the other hand, while a larger beam size exhibits less divergence, it becomes more susceptible to turbulence. Therefore, achieving an optimal beam size is crucial for improved performance. The gain of the transmitter beam [26], $G_{t}$ is given in (3).
${\Theta_{B}}$ is related to wavelength $\lambda$ and beam size $W_{0}$ as $\Theta_{B}=\frac{\lambda}{\pi W_{0}}$.
@@ 2.2 Path loss:
Given its proximity to Earth, LEO experiences minimal diffraction loss, making it particularly advantageous for quantum key distribution (QKD) experiments. However, orbits with slower speeds, such as the Middle Earth Orbit (MEO) or GEO, offer the benefit of maintaining a continuous link for extended periods, allowing for prolonged QKD operations. Nonetheless, it’s worth noting that MEO and GEO orbits come with their own challenges, including higher radiation levels and increased propagation loss.
Free space propagation loss [30] in dB is $\log_{10}L_{r}$ where $L_{r}$ is given in (3).
@@ 2.2.1 Atmospheric loss:
$\eta_{atte}$ is due to the static components of the atmosphere and $\eta_{tur}$ [31]is due to the turbulence i.e dynamic components.
The Static components constitute various types of gases, water vapors, and dust particles that constitute the static atmosphere. These atmospheric elements absorb energy from the photons and this phenomenon is known as absorption. It depends on the wavelength. Photon collisions with air particles cause scattering. The dynamic components are discussed in the next subsection 2.2.2.
@@
Table 2: Atmospheric parameters considered for MODTRAN simulations for IAO Hanle.
Figure 1: Simulated atmospheric transmittance at the IAO Hanle, a rural location. The X-axis represents the wavelength in nm, ranging from 500 to 1500, while the Y-axis represents the transmittance.
@@ 2.2.2 Turbulence
[…]
@@
One can also compute $r_{0}$ using the ‘seeing parameter’. We have used this method to calculate $r_{0}$ for different Indian ground station locations [36]. When there is no turbulence present, the parameter $r_{0}$ represents the diameter of an equivalent aperture where the resolution of a telescope is approximately diffraction-limited. A larger value of $r_{0}$ implies that turbulence has a lesser impact on the propagation of the beam and vice versa. Since $r_{0}$ varies with the wavelength according to $\lambda^{6/5}$, longer operating wavelengths result in larger values of the $r_{0}$. This suggests that at longer wavelengths, the effect of turbulence on the wavefront is less severe.
One more way to express the strength of turbulence is $\frac{D}{r_{0}}$. This is a ratio of the diameter of the beam from the telescope aperture to the atmosphere’s coherence length. The higher the value, the more turbulated the beam. As the value for $\frac{D}{r_{0}}$ increases higher order wavefront distortion will accumulate  [37]. Hence objective is to have minimum $\frac{D}{r_{0}}$. But its value shouldn’t be <1<1< 1. Though no wavefront distortion will be there but optical beam will go through a beam wander phenomenon which makes it difficult to detect the beam at the receiver.
@@ 3 Link budget calculation for beacon beams
To establish and sustain a free-space optical link, it is essential for the spacecraft’s altitude determination and control system to ensure precise pointing accuracy toward the optical ground station. Conversely, a laser beacon tracking system has the capability to furnish accurate ground-based altitude information, thus facilitating communication from the satellite. A laser beacon is transmitted from the optical ground station with a different wavelength from the quantum signal beam. The satellite will also transmit a beacon beam to get the position of the ground station. For the uplink, we use 532 nm, and for the downlink beacon 1550 nm wavelength. For both uplink and downlink beacons, we fix the beam divergence to be $500\mu rad$. This high divergence leads to footprints of the beacon lasers for the best possible tracking at the receiving ends. Another key advantage of choosing high-divergence beacon beams is the low pointing error loss during transmission. The link budget calculation for the beacon beam is also given in Table 3 for a ground telescope of 15 cm aperture and a satellite telescope of 30 cm aperture,[43]. The calculation for the optimum size of 15 cm is given in section A.4 of the Appendix.
No Parameter unit Uplink Downlink
1 Tx power ($P_{t}$) W 1.0 1.0
2 Wavelength ($\lambda$) nm 532 1550
3 Tx Beam diameter mm 150.0 300.0
4 Tx Beam divergence ($2\Theta_{B}$) $\mu rad$ 500 500
5 Tx gain ($G_{t}$) dB 81.07 81.07
6 Tx optics loss ($\eta_{t}$) dB -2.20 -2.20
7 Propagation distance ($L$) km 500 500
8 Path loss ($L_{r}$) dB -261.48 -252.16
9 Atmospheric transmittance ($\eta_{atm}$)=0.73/0.81 dB -1.36 -0.9
10 Turbulence ($\eta_{tur}$)=0.64/0.96 dB -1.88 -0.18
11 RX antenna diameter cm 30.0 15.0
12 RX gain ($G_{r}$) dB 124.97 109.66
13 RX optics loss ($\eta_{r}$) dB -2.2 -2.2
14 Total loss dB 63.08 66.91
Table 3: Link budget for uplink 532 nm and downlink 1550 nm beam. Rows 1 to 6 represent parameter for the transmitter telescope. Rows 7 to 10 represent propagation losses. Rows 11 to 13 represent parameters for the receiver telescope.
To summarize the information contained in the various tables, the first Table 1 provides the link budget for the signal beam at the IAO Hanle observatory. The atmospheric parameters for the period from January to December 2015 which are sourced from the GIOVANNI database presented in Table 2. Table 3 presents the link budget for the uplink and downlink beacon beams at IAO Hanle. Finally, Table 4 compares the link budget analysis for the signal and beacon beams between IAO Hanle, Aries Nainital, and Mount Abu. For a comprehensive understanding of the parameters employed to compute the link budget for these two observatories, please refer to Table A4 within the Appendix section. The beacon beam at 532 nm is used for uplink and 1550 nm for downlink, whereas the last column for the signal beam is at 810 nm. For the downlink beacon, due to the smaller size of the ground telescope, the loss is always greater compared to the uplink.
Location Uplink beacon loss in dB Downlink beacon loss in dB Signal loss in dB
IAO Hanle 63.08 66.91 35.91
ARIES Nainital 67.22 67.54 37.78
Mount-Abu 66.26 66.93 37.19
Table 4: Link budget analysis for 3 different locations in India for uplink beacon, downlink beacon and signal beam of wavelength 532 nm, 1550 nm and 810 nm respectively.
@@ 7 Propagation distance ($L$) km 500 500
When the zenith angle is zero degrees, the loss is indicated in Table 4 for the three separate places in India. The air mass will increase when the satellite moves at a greater zenith angle, which is the cause of an increase in loss in dB, as depicted in Fig.2. The plot is given for one of the locations i.e. IAO Hanle.
Figure 2: Atmospheric transmission at different zenith angles. X-axis represents the zenith angles in degree. Y-axis represents the loss in dB which will accumulate with increase in angle.
@@
3.1 Doppler effect assessment
In the case of a LEO satellite, the apparent velocity relative to the ground exhibits a linear increase as the satellite’s elevation angle rises. This scenario involves an analysis of the Doppler effect. As per the Doppler effect, the central frequency of the signal undergoes substantial changes during the data transmission and communication process, primarily due to the rapid motion of LEO satellites. If the satellite receiving end moves away from the transmitting ground end, the signal frequency will decrease, potentially resulting in a higher error rate upon signal reception and recovery. Normalized Doppler $\frac{\Delta f}{f}$ is given by  [44]
$\frac{\Delta f}{f}=\frac{-1}{c}\frac{r_{E}r\sin(\psi(t)-\psi(t_{0}))\cos(\cos^{-1}(\frac{r_{E}}{r}\cos\theta_{max})-\theta_{max})\omega_{f}(t)}{\sqrt{r_{E}^{2}+r^{2}-2r_{E}r\sin(\psi(t)-\psi(t_{0}))\cos(\cos^{-1}(\frac{r_{E}}{r}\cos\theta_{max})-\theta_{max})}}$ (20)
where,
  • $r_{E}$= radius of earth which is 6400 km
  • $r$ = position of satellite w.r.t center of earth which is 6900 km
  • $\psi(t)-\psi(t_{0})$ is satellite visibility window duration which is 8 min. approx.
  • $\theta_{max}$ is maximum elevation angle which is 909090 degrees.
  • $\omega_{f}(t)$ angular speed of satellite which is 0.00105 rad/sec
  • $c$ is the velocity of light.
For 810 nm signal beam the $\frac{\Delta f}{f}$ comes out to be $1.5\times 10^{-5}$. Correspondingly the frequency shift $\Delta f$ w.r.t the center carrier frequency 380 THz is 0.0057 THz. This shift is considered negligible.
4 Polarization-based quantum communication protocol simulations
Polarization is a fundamental property of photons given by the plane in which the electric field vector oscillates. Polarization states can be represented as vectors in a two-dimensional Hilbert space[45]. Different mutually unbiased bases in this degree of freedom, like {horizontal (H), vertical (V)}, {diagonal (D) and anti-diagonal (A)} and {right-circular (R) and left-circular (L)}, can be used to create bipartite polarization-entangled states. For example, an entangled pair can be represented as
$\psi=\frac{1}{\sqrt{2}}(H_{A}V_{B}+V_{A}H_{B})$ (21)
where subscripts A and B denote two different photons.
Polarization states are used to encode information. To measure the polarization of a photon, polarizing beam splitters and waveplates are used. The measurement outcomes depend on the chosen basis (e.g., H/V basis or D/A basis). In the following section 4.1, we discuss atmospheric effects on polarization, that can cause error in measurement results. Protocols like BB84 with weak coherent pulse sources use polarization states to generate and share cryptographic keys. Details are given in Sec 4.2, whereas QKD using an Entangled source is discussed in Sec 4.3.
4.1 Effect of atmosphere on polarization degree of freedom
In Free-Space QKD, one of the key challenges is the impact of atmospheric turbulence on the polarization state of photons used for encoding quantum information. As polarized photons propagate through the atmosphere, they are subject to turbulence, which can cause random fluctuations in the refractive index of the air causing an effect of birefringence (See the discussion in 2.2.2). These fluctuations affect the phase of orthogonal polarization components differently, leading to a change in overall polarization state of the photons. This leads to a higher Quantum Bit Error Rate (QBER), as the transmitted polarization state may not match the intended one. These effects of turbulence on polarization state and possible mitigation strategies are as follows:
Depolarization: Turbulence can mix different polarization components, converting pure polarized states into partially or completely depolarized mixed states. Such errors in the real world implementation can be countered using adaptive optics based solution that corrects for any wavefront distortion caused during transmission.
Birefringence: Variations in refractive index along different propagation axes create an effect of birefringence, causing phase shifts between different polarization components. There have been successful demonstration of various active[46, 47] and passive [48] polarization correction techniques that corrects for any polarization scrambling because of birefringence in free-space or in fiber based systems.